OpenStudy (anonymous):

17c4 simplify the expression

OpenStudy (lgbasallote):

\[17C4 \implies \frac{17}{4!(17-4)!}\]

OpenStudy (lgbasallote):

\[\implies \frac{17!}{4!13!}\] can you do it from here now?

OpenStudy (anonymous):

i dont even get how to do this?

OpenStudy (lgbasallote):

\[nCr \implies \frac{n!}{r!(n-r)!}\] are you aware of this formula?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

do you have notes on this topic?

OpenStudy (anonymous):

no